# Federated attention contrastive learning models for prostate cancer diagnosis and Gleason grading

### v4

@@ Abstract
Artificial intelligence (AI) shows great promise in revolutionizing medical imaging, improving diagnosis, and refining treatment methods. However, the training of AI models relies on extensive multi-center datasets, presenting a potential challenge due to concerns about data privacy protection. Federated learning offers a solution by enabling a collaborative model across multiple centers without sharing raw data. In this study, we present a Federated Attention Contrastive Learning (FACL) framework designed to address challenges associated with large-scale pathological images and data heterogeneity. FACL improves model generalization by maximizing attention consistency between the local client and the server model. To enhance privacy and validate robustness, we incorporate differential privacy by introducing noise during parameter transfer. We assess the effectiveness of FACL in cancer diagnosis and Gleason grading tasks using 19,461 whole slide images of prostate cancer sourced from multiple centers. In the diagnosis task, FACL achieves an area under the curve (AUC) of 0.9718, outperforming seven centers whose average AUC is 0.9499 when categories are relatively balanced. In the Gleason grading task, FACL attained a Kappa score of 0.8463, surpassing the average Kappa score of 0.7379 from six centers. In conclusion, FACL offers a robust, accurate, and cost-effective AI training model for prostate cancer pathology while maintaining effective data safeguards.
@@ 1 Introduction
Prostate cancer ranks as the second most prevalent cancer and the fifth leading cause of cancer-related deaths among men sung2021global . The gold standard for solid tumor diagnostics involves examining hematoxylin and eosin (H&E) stained whole slide images (WSIs) lu2015automated ; schirris2022deepsmile ; chikontwe2022feature . Accurate diagnosis and Gleason grading of prostate cancer WSIs are essential for effective clinical management and treatment kartasalo2021artificial ; ma2023identification . Specifically, Gleason grading is the most reliable method to assess prostate cancer aggressiveness delahunt2012gleason ; silva2020going ; short2019gleason . However, interobserver and intraobserver variability in Gleason scores can lead to under or over-treatment of patients in real-world scenarios kartasalo2021artificial ; GEORGE2022262 . Furthermore, manual diagnostic labeling of pathological images is time-consuming and costly campanella2019clinical ; wang2022sclwc ; wang2023retccl ; xiang2023automatic .
@@
Recent advancements in artificial intelligence (AI) have propelled its expanded utilization in pathological image diagnosis. Pioneering research studies have showcased the capability of AI to provide precise, cost-effective, and scalable solutions by automating the analysis of patient data kartasalo2021artificial ; bulten2022artificial ; bulten2020automated ; takahashi2023artificial . Numerous previous studies campanella2019clinical ; bulten2020automated ; perincheri2021independent ; nagpal2020development ; pantanowitz2020artificial ; mun2021yet have unequivocally validated AI models’ effectiveness in diagnosing prostate cancer and performing Gleason grading tasks. However, it is important to note that these AI models necessitate a substantial amount of data for training to attain optimal performance muti2021development ; echle2020clinical ; saldanha2022swarm . As a result, a typical strategy involves the collection of data from multiple centers, including medical centers and hospitals, and subsequently training the AI models through a centralized learning process. Nevertheless, it is crucial to acknowledge that the transfer of extensive volumes of sensitive patient data among clinics, hospitals, and other medical institutions can potentially compromise patient privacy, and may run counter to data protection regulations. Moreover, there are notable technical challenges to overcome, such as ensuring efficient data transfer and addressing storage requirements, which present considerable hurdles in this endeavor saldanha2022swarm ; pati2022federated .
@@
Federated learning (FL) represents an innovative paradigm in which models are trained by sharing model parameter updates from decentralized data sources pati2022federated ; liu2021federated ; li2020federated ; xu2021federated ; sheller2020federated . Retaining data locally can effectively address issues related to excessive data transmission and data ownership rieke2020future . By directly sharing model parameters rather than the private data of multiple centers, FL offers a promising solution for applications involving sensitive data. In general, FL training methods can be categorized into two types: (1) primary server and (2) peer-to-peer lu2022federated . The key distinction lies in whether a primary server is employed to collect each client’s model parameters. FL has demonstrated significant potential for sensitive data analytics across multiple healthcare domains xu2021federated . Examples include electronic health records (EHR) brisimi2018federated ; boughorbel2019federated ; huang2019patient ; min2019predictive ; duan2020learning , healthcare IoT (Internet of Things) chen2020fedhealth ; brophy2021estimation , drug discovery choudhury2019predicting ; chen2021fl ; xiong2022facing , and medical image analysis (e.g., X-ray, Ultrasound, CT, MRI, PET, WSIs) yan2021experiments ; feki2021federated ; lee2021federated ; dou2021federated ; florescu2022federated ; guo2021multi ; stripelis2021scaling ; shiri2021federated ; andreux2020siloed ; adnan2022federated . Despite the advantages of FL in the context of gigapixel WSIs andreux2020siloed ; wang2023generalizable ; baid2022federated ; ogier2023federated , significant challenges remain, particularly in ensuring robustness in the face of considerable heterogeneity among WSIs with a wide range of morphological variations.
@@
Data heterogeneity across different centers, characterized by each local model possessing non-identically and independently distributed (Non-IID) data, presents a significant challenge in the field of FL kairouz2021advances . The divergence between the training and test data distributions can substantially impact model robustness negatively de2022mitigating . Several studies have made notable progress in addressing the heterogeneity problem through the following approaches: (1) Variance reduction: Recent advancements in this area focus on balancing the data distribution obtained by centers from other centers ahn2022communication , to reduce variance. This is achieved by considering various aspects such as parameters, update directions, or feature representations among local models nguyen2022federated . (2) Global model update: To expedite model convergence and enhance robustness, adaptive optimization mcmahan2017communication ; li2021fedrs ; gao2022feddc or personalized optimizationhanzely2021personalized techniques are employed during the model parameter aggregation process. However, it should be noted that these improvements do not specifically address the heterogeneity issue of histopathological images in FL. By employing these approaches, researchers have made significant strides in mitigating data heterogeneity challenges in FL. Nevertheless, additional research is required to address the unique characteristics and complexities associated with histopathological images in this context.
@@
The accurate diagnosis of pathological images requires a detailed analysis of the distinctive features present within cancerous regions. In this study, we have integrated an attention mechanism into the foundational classification model to facilitate this meticulous examination. Moreover, to effectively mitigate the inherent heterogeneity in federated learning settings, we have introduced contrastive learning techniques into the attention distribution process of the model. Consequently, we propose a novel framework termed Federated Attention Contrastive Learning (FACL), specifically devised to bolster attention consistency between the local client models and the central server model. This augmentation is strategically aimed at enhancing the generalization capacity of the model, particularly towards unseen data instances. Moreover, while FL retains data locally, there is an inherent risk of private information leakage during the parameter transfer to the server wu2022adaptive ; wang2022safeguarding . To address this concern, we incorporate the use of differential privacy dwork2006our as a protective measure zhao2022cork . It is worth noting that most existing advancements in this field have been proposed and validated using a limited number of small-scale image datasets, such as MNIST, CIFAR-10, CIFAR-100, EMNISTL, and Tiny ImageNet nguyen2022federated ; li2021fedrs ; li2021model . To validate the effectiveness of our proposed model, we have curated a large-scale clinical dataset consisting of 19,461 WSIs obtained from multiple centers. Rigorous testing conducted on this dataset, specifically focusing on prostate cancer diagnosis and Gleason grading, demonstrates clinical-grade performance.
Our research makes significant contributions in the following areas:
(1) We proposed a novel FL framework named FACL, which ensures privacy preservation and minimizes unnecessary data transmission, aiming at achieving better diagnosis and Gleason grading of prostate cancer based on large-scale and multi-center pathology images.
(2) We introduced a contrastive attention algorithm to address heterogeneity among FL centers. This algorithm improves the model’s generalization ability by maximizing attention consistency between local client and server models.
(3) To comprehensively evaluate our approach, we curated a large-scale clinical prostate cancer dataset comprising 19,461 WSIs from multiple medical centers. These diverse data distributions ensure that our evaluation captures the complexities inherent in real-world prostate cancer cases.
@@ 2 Related works
This section presents a comprehensive overview of federated and contrastive learning, with a specific emphasis on their applications in medical image analysis.
@@ 2.1 Federated learning
Following Google’s introduction of the FL framework FedAvg mcmahan2017communication , it has undergone extensive development, particularly addressing challenging aspects such as non-independent and identically distributed and unbalanced properties.
To address the Non-IID problem, FedDif ahn2022communication introduced a diffusion mechanism where local models diffuse in the network, enabling them to learn data distributions beyond their initial experience before global aggregation. Following sufficient diffusion iterations, each local model accumulated personalized data distributions, resembling the effects seen with training identically and independently distributed data. ComFed nguyen2022federated employed variance reduction techniques on the client side to streamline server aggregation, coupled with global adaptive update techniques for accelerated learning. FedRS li2021fedrs proposed the use of “Restricted Softmax” to control the update of missing class weights during the local process. FedDC gao2022feddc introduced a lightweight modification in the local training phase, employing an auxiliary local drift variable for tracking the gap between local and global model parameters. FedProx litian2020federated , as an improvement upon FedAvg, incorporated a proximal term into the local objective function. This addition significantly enhanced the stability of the method, showcasing its effectiveness. FedFAME malaviya2023fedfame introduces a comprehensive framework centered on model contrastive learning, with a primary focus on tackling challenges related to local client data enhancement and Non-IID issues, particularly in domains like text and graphs. While these methods have significantly contributed to Non-IID problem-solving, it’s crucial to acknowledge their primary validation on conventional natural images. Consequently, these approaches might not seamlessly apply to the distinctive characteristics of large-scale pathological images prevalent in medical imaging datasets.
FL has garnered widespread recognition for its effectiveness and efficiency in healthcare applications. Dou et al. dou2021federated exemplified the feasibility of FL in detecting COVID-19-related CT abnormalities through a multinational study, achieving external validation and demonstrating robust generalization on internal and external datasets. In another study, Guo et al. guo2021multi introduced a cross-site MR image reconstruction model that aligns learned intermediate latent features between different source centers with the latent feature distribution at the target centers. This innovative approach contributes to enhancing the model’s performance and adaptability across diverse datasets. Pati et al. pati2022federated applied FL to rare disease diagnosis, leveraging data from 71 centers across six continents. The outcome was an automated tumor boundary detector for rare glioblastoma, showcasing the efficacy and practicality of FL in a multicenter setting. Liu et al. liu2023fedcl took a novel approach by combining FL and contrastive learning in experiments with small-sized medical images, introducing FedCL to address data heterogeneity challenges. It is crucial to note that FedCL’s multicenter approach involves dividing a dataset into multiple subsets, essentially creating a pseudo multicenter. However, this algorithm faces limitations in performing feature extraction in every round on large-sized pathological images. These studies collectively underscore the potential of FL in revolutionizing healthcare, emphasizing its adaptability to diverse scenarios while acknowledging certain constraints in specific applications.
@@
FedAvg mcmahan2017communication Natural images MNIST CIFAR-10 Shakespeare
FedDif ahn2022communication Natural images MNIST FMNIST CIFAR-10
ComFed nguyen2022federated Natural images CIFAR-10
FedRS li2021fedrs Natural images CIFAR-10
FedDC gao2022feddc Natural images CIFAR-10
FedProx litian2020federated Natural images MNIST Sent140 FEMNIST Shakespeare
@@
FedFAME malaviya2023fedfame Natural images CIFAR-10 20NewsGroups Fashion-MNIST 5AbstractsGroup
FedCL liu2023fedcl Medical images (X-ray) 5908 images
dou2021federated Medical images (CT) 10594 images
guo2021multi Medical images (MR) 3443 images
pati2022federated Medical images (ET, TC, WT) 6314 images
saldanha2022swarm Pathological images Epi700 (661) TCGA (632) DACHS (2448) QuASAR (2190) YCR BCIP (889)
lu2022federated Pathological images RCC (1184) BRCA (2126) CCRCC (511)
Prop-FFLhosseini2023proportionally Pathological images Kidney (424) Lung (1217)
@@
Several studies have showcased the potential of FL in analyzing histopathological images. Saldanha et al. saldanha2022swarm employed the FL model to predict BRAF mutation status and microsatellite instability from H&E stained colorectal cancer pathology sections. Ming et al. lu2022federated and Seyedeh et al. hosseini2022cluster both utilized FL and differential privacy to underscore its efficacy in safeguarding medical privacy data on the public dataset TCGA. Addressing the issue of fairness among participants, Hosseini et al. hosseini2023proportionally introduced proportional fair federated learning (Prop-FFL) and substantiated its effectiveness on two pathological datasets. As shown in Table 1, FL finds predominant application in the realm of natural images. However, in the field of medical imaging, its usage is more common for non-pathological images. Notably, when it comes to pathological images, public datasets are predominantly utilized. However, a prevalent limitation in FL research applied to histopathological images lies in the lack of data diversity. Many studies rely heavily on public datasets, thereby lacking comprehensive verification of actual clinical data.
@@
Fig. 1: The framework of feature extraction and model training process for FACL model on pathological images.
2.2 Contrastive learning
Contrastive learning (CL) stands out as a potent paradigm for visual representation learning, captivating considerable research attention in recent years. Seminal frameworks like SimCLR chen2020simple and MoCo he2020momentum have paved the way for this method, sparking a wave of subsequent developments. At its essence, contrastive learning revolves around discrimination minimizing the distance between augmented view representations of the same image (termed positive pairs) while maximizing the distance between augmented views of different images (referred to as negative pairs) li2021model ; krishnan2022self . This fundamental principle underlies the remarkable progress in contrastive learning.
In the healthcare domain, prior research has showcased performance enhancements with the application of contrastive learning in specific domains such as electronic health records (EHR) wanyan2021contrastive ; li2022multi , human activity recognition (HAR) tang2020exploring , electrocardiography (ECG) diamant2022patient and medical image analysis vu2021medaug ; chaitanya2020contrastive ; azizi2021big ; li2021dual ; wang2023retccl .
In histopathology image analysis, several studies have explored the application of contrastive learning to enhance performance. Wang et al. wang2023retccl devised a CL-based backbone model that integrates multiple sub-memories and incorporates population-level discrimination to diminish potential false negatives. They trained this model under the conventional contrastive learning assumption. Yan et al. yan2022deep utilized a deep clustering algorithm to maximize similarity between instance and cluster projection representations through contrastive learning. Simultaneously, they learned semantic features and generated initial clusters from unlabeled WSI patches. Zhu et al. zhu2022murcl introduced a novel framework, multi-instance reinforced contrastive learning (MuRCL), to deeply mine intrinsic semantic relations among different patches, thereby advancing WSIs classification. Fashi et al. fashi2022self proposed a fully supervised contrastive learning setting to enhance representation robustness for WSIs classification and perform pretext searches and downstream tasks. However, these contrastive learning methods primarily focus on feature representation and have not yet been tested on large-scale, diverse histopathology image datasets. Therefore, we extend this concept beyond visual representation to attention contrastive model training.

### v5

# Federated attention consistent learning models for prostate cancer diagnosis and Gleason grading

@@ Abstract
Artificial intelligence (AI) holds significant promise in transforming medical imaging, enhancing diagnostics, and refining treatment strategies. However, the reliance on extensive multicenter datasets for training AI models poses challenges due to privacy concerns. Federated learning provides a solution by facilitating collaborative model training across multiple centers without sharing raw data. This study introduces a federated attention-consistent learning (FACL) framework to address challenges associated with large-scale pathological images and data heterogeneity. FACL enhances model generalization by maximizing attention consistency between local clients and the server model. To ensure privacy and validate robustness, we incorporated differential privacy by introducing noise during parameter transfer. We assessed the effectiveness of FACL in cancer diagnosis and Gleason grading tasks using 19,461 whole-slide images of prostate cancer from multiple centers. In the diagnosis task, FACL achieved an area under the curve (AUC) of 0.9718, outperforming seven centers with an average AUC of 0.9499 when categories are relatively balanced. For the Gleason grading task, FACL attained a Kappa score of 0.8463, surpassing the average Kappa score of 0.7379 from six centers. In conclusion, FACL offers a robust, accurate, and cost-effective AI training model for prostate cancer pathology while maintaining effective data safeguards.
@@ 1 Introduction
Prostate cancer ranks as the second most prevalent cancer and the fifth leading cause of cancer-related deaths among men [1]. The standard approach for diagnosing solid tumors involves examining hematoxylin and eosin (H&E)–stained whole-slide images (WSIs) [2, 3, 4]. Accurate diagnosis and Gleason grading of prostate cancer WSIs are crucial for effective clinical management and treatment [5, 6]. Gleason grading, in particular, is the most reliable method for assessing aggressiveness [7, 8, 9]. However, interobserver and intraobserver variability in Gleason scores can result in under or over-treatment of patients in real-world scenarios [5, 10]. Moreover, manually labeling pathological images for diagnosis is time-consuming and expensive [11, 12, 13, 14].
@@
Recent advancements in artificial intelligence (AI) have accelerated its application in pathological image diagnosis. Groundbreaking research studies have demonstrated AI’s ability to deliver precise, cost-effective, and scalable solutions by automating patient data analysis [5, 15, 16, 17]. Numerous prior studies have unequivocally confirmed the effectiveness of AI models in diagnosing prostate cancer and conducting Gleason grading tasks [11, 16, 18, 19, 20, 21]. However, it is essential to recognize that these AI models demand a significant volume of training data to achieve optimal performance [22, 23, 24]. Consequently, a common approach involves gathering data from various centers, including medical institutions and hospitals, and then training AI models through centralized learning processes. Nevertheless, it is critical to acknowledge that transferring extensive amounts of sensitive patient data among clinics, hospitals, and other medical establishments may jeopardize patient privacy and potentially conflict with data protection regulations. Additionally, there are notable technical challenges to surmount, such as ensuring efficient data transfer and addressing storage requirements, which pose significant obstacles in this pursuit [24, 25].
@@
Federated learning (FL) presents an innovative paradigm wherein models are trained by sharing model parameter updates from decentralized data sources [25, 26, 27, 28, 29]. Local retention of data effectively addresses issues related to excessive data transmission and ownership [30]. By directly sharing model parameters rather than private data from multiple centers, FL offers a promising solution for applications involving sensitive data. FL training methods generally fall into two categories: (1) primary server and (2) peer-to-peer [31], with the key distinction being whether a primary server is employed to collect model parameters for each client. FL has demonstrated significant potential for sensitive data analytics across various healthcare domains [28], including electronic health records (EHR) [32, 33, 34, 35, 36], healthcare Internet of Things (IoT) [37, 38], drug discovery [39, 40, 41], and medical image analysis (e.g., X-ray, Ultrasound, CT, MRI, PET, WSIs) [42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Despite the advantages of FL in gigapixel WSIs [50, 52, 53, 54], significant challenges remain, particularly in ensuring robustness against considerable heterogeneity among WSIs with a wide range of morphological variations
@@
Data heterogeneity across different centers, characterized by each local model possessing non-identically and independently distributed (non-IID) data, presents a significant challenge in FL [55]. The divergence between the training and test data distributions can substantially negatively impact model robustness [56]. Several studies have made notable progress in addressing the heterogeneity problem through the following approaches. (1) Variance reduction: Recent advancements in this area have focused on balancing the data distribution obtained by centers from other centers to reduce variance [57]. This is achieved by considering various aspects such as parameters, update directions, or feature representations among local models [58]. (2) Global model update: To expedite model convergence and enhance robustness, adaptive [59, 60, 61] and personalized optimization [62] techniques were employed during the model parameter aggregation process. However, it should be noted that these improvements do not specifically address the heterogeneity issue of histopathological images in FL. By employing these approaches, researchers have made significant strides in mitigating data heterogeneity challenges in FL. Nevertheless, additional studies are required to address the unique characteristics and complexities of histopathological imaging in this context.
Accurate diagnosis of pathological images requires a detailed analysis of features within cancerous regions. In this study, we employed a multiple-instance learning classification model with an attention mechanism to facilitate meticulous examination. From a practical perspective, the model identifies the region of interest of the pathology image and the necessary diagnostic results for physicians, indicating potential applications in pathological imaging. To enhance the final performance of each client model within the FL framework, we aimed to improve the alignment of attention distribution between each client and its corresponding server counterpart models. To achieve this goal, we introduced a novel framework called federated attention consistent learning (FACL). FACL involves two iterative steps: individual client model training and server model fusion. Initially, each client model underwent training using its private dataset. Subsequently, the learned weights from these client models were aggregated to update the central server model. To ensure that each client model accurately reflected the capabilities of the server model, we replicated the server model as an additional replica model for each client in the initial step. Furthermore, to enhance the performance of each client model, we enforced an attentional consistency constraint, ensuring that each client model remained consistent with the server counterpart model when processing the same input images. This enhancement strategy aims to bolster the model’s generalization capability, particularly for data samples that have not been encountered previously.
@@
Moreover, although FL retains data locally, there is an inherent risk of private information leakage during parameter transfer to the server [63, 64]. To address this concern, we incorporated differential privacy [65] as a protective measure [66]. Notably, most existing advancements in this field have been proposed and validated using a limited number of small-scale image datasets, such as MNIST, CIFAR-10, CIFAR-100, EMNISTL, and Tiny ImageNet [58, 60, 67]. To validate the effectiveness of the proposed model, we curated a large-scale clinical dataset consisting of 19,461 WSIs obtained from multiple centers. Rigorous testing conducted on this dataset, specifically focusing on prostate cancer diagnosis and Gleason grading, demonstrated the clinical grading performance.
Our research makes significant contributions in the following areas: (1) We propose a novel FL framework named FACL, which ensures privacy preservation and minimizes unnecessary data transmission, aiming to achieve better diagnosis and Gleason grading of prostate cancer based on large-scale and multicenter pathology images. (2) We introduced an attention consistency algorithm to address the heterogeneity among FL centers. This algorithm improves the generalization ability of the model by maximizing the attention consistency between the local client and server models. (3) To comprehensively evaluate our approach, we curated a large-scale clinical prostate cancer dataset comprising 19,461 WSIs from multiple medical centers. These diverse data distributions ensure that our evaluation captures the complexities inherent in real-world prostate cancer cases.
@@ 2 Related works
This section presents a comprehensive overview of federated learning, specifically emphasizing its applications in medical image analysis.
@@ 2.1 Federated learning
Following Google’s introduction of the FL framework FedAvg [59], it has developed extensively, particularly in addressing challenging aspects such as non-IID and unbalanced properties.
@@
FedAvg [59] Natural images MNIST CIFAR-10 Shakespeare
FedDif [57] Natural images MNIST FMNIST CIFAR-10
ComFed [58] Natural images CIFAR-10
FedRS [60] Natural images CIFAR-10
FedDC [61] Natural images CIFAR-10
FedProx [68] Natural images MNIST Sent140 FEMNIST Shakespeare
@@
FedFAME [69] Natural images CIFAR-10 20NewsGroups Fashion-MNIST 5AbstractsGroup
[45] Medical images (CT) 10594 images
[47] Medical images (MR) 3443 images
[25] Medical images (ET, TC, WT) 6314 images
[24] Pathological images Epi700 (661) TCGA (632) DACHS (2448) QuASAR (2190) YCR BCIP (889)
[31] Pathological images RCC (1184) BRCA (2126) CCRCC (511)
Prop-FFL[70] Pathological images Kidney (424) Lung (1217)
To address the non-IID problem, FedDif [57] introduced a diffusion mechanism in which local models diffused in a network, enabling them to learn data distributions beyond their initial experience prior to global aggregation. Following sufficient diffusion iterations, each local model accumulated personalized data distributions resembling the effects observed with training using identically and independently distributed data. ComFed [58] employs variance reduction techniques on the client side to streamline server aggregation, coupled with global adaptive update techniques for accelerated learning. FedRS [60] proposed the use of ”Restricted Softmax” to control the update of missing class weights during the local process. The FedDC [61] introduced a lightweight modification in the local training phase by employing an auxiliary local drift variable to track the gap between the local and global model parameters. As an improvement on FedAvg, FedProx [68] incorporated a proximal term into the local objective function. This significantly enhanced the stability of the method, demonstrating its effectiveness. Further, FedFAME [69] introduced a comprehensive framework centered on model-contrastive learning with a primary focus on tackling challenges related to local client data enhancement and non-IID issues, particularly in domains such as texts and graphs. Although these methods have significantly contributed to non-IID problem-solving, it is crucial to acknowledge their primary validation on conventional natural images. Consequently, these approaches cannot be seamlessly applied to the distinctive characteristics of large-scale pathological images prevalent in medical imaging datasets.
FL has garnered widespread attention because of its effectiveness and efficiency in healthcare applications. Dou et al. [45] demonstrated the feasibility of FL in detecting COVID-19-related CT abnormalities through a multinational study, achieving external validation and demonstrating robust generalization on internal and external datasets. In another study, Guo et al. [47]introduced a cross-site MR image reconstruction model that aligned the learned intermediate latent features between different source centers with the latent feature distribution at the target centers. This innovative approach enhanced the performance and adaptability of the model across diverse datasets. Pati et al. [25] applied FL to rare disease diagnosis, leveraging data from 71 centers across six continents. The outcome was an automated tumor boundary detector for rare glioblastomas, demonstrating the efficacy and practicality of FL in multicenter settings. These studies collectively underscore the potential of FL in revolutionizing healthcare, emphasizing its adaptability to diverse scenarios, while acknowledging certain constraints in specific applications.
@@
Several studies have demonstrated the potential of FL for analyzing histopathological images. Saldanha et al. [24] employed the FL model to predict the BRAF mutation status and microsatellite instability in H&E-stained colorectal cancer pathology sections. Ming et al. [31] and Seyedeh et al. [71] utilized FL and differential privacy to underscore their efficacy in safeguarding medical privacy data in the public TCGA dataset. Addressing the issue of fairness among participants, Hosseini et al. [70] introduced proportional fair federated learning (Prop-FFL) and substantiated its effectiveness on two pathological datasets. As illustrated in Table 1, FL finds predominant applications in natural images. However, in the field of medical imaging, its use is more common for non-pathological images. Public datasets are predominantly used for pathological imaging. However, a prevalent limitation of FL research applied to histopathological images is the lack of data diversity. Several studies rely heavily on public datasets, thereby lacking a comprehensive verification of actual clinical data.
@@
Fig. 1: Framework of feature extraction and model training process for the FACL model on pathological images